Grade upto college level Wave Motion

A parallel beam of white light is incident normally on a water film 1.0 × 10-4 cm thick. Find the wavelength in the visible range (400 nm-700 nm) which are strongly transmitted by the film. Refractive index of water = 1.33.

Amit Saxena
12 Years agoGrade upto college level

1 Answer

Navjyot Kalra
12 Years ago
Sol. Given d = 0.0011 * 10^–3 m For minimum reflection of light, 2μd = nλ| ⇒ μ = nλ/2d = 2nλ/4d = 580 * 10^-9 * 2n/4 * 11 * 10^-7 = 5.8/44(2n) = 0.132 (2n) Given that, μ has a value in between 1.2 and 12. and 1.5. ⇒ When, n = 5 μ = 0.132 = 10 = 1.32.